y=x-3
y=-x+9
What is the solution?

Answer by stanbon(75887)

y=x-3
y=-x+9
----
x-3 = -x+9
2x = 12
x = 6
---
Solve for "y":
y = x-3
y = 3
----
solution: (6,3)
==================
Cheers,
Stan H.
===================